Given the thermodynamic data below, calculate the value of the equilibrium constant for the reaction shown at 25.0ºC

H₂ (g) + I₂ (g) ⇄ 2 HI (g)

ΔHº = -9.48 kJ

ΔSº = +21.79 J/K

K = Answer at 25.0ºC
Answer #1

AG =ón - TDS 04:19.48 - (25+273 ) KX 21.793/k az - 9:48 x 10² ] - 6493.42] Da= -15973-423 su= - RT enk A lnk= Da - RT lnk= -